OpenStudy (anonymous):

Find dy / dx by implicit differentiation. xy + 5x + 3x^2 = 5... I got y'=(-y-5-6x)/x...but i got stuck

OpenStudy (amistre64):

[xy]' = x'y + xy' [5x]' = 5 [3x^2]'=6x [5]' = 0

OpenStudy (amistre64):

x'y + xy' +5 + 6x = 0 y' = (-5-6x -y)/x

OpenStudy (anonymous):

yes...i got that steps i took all of the derviatives

OpenStudy (amistre64):

where you stuck at? you found the derivative

OpenStudy (anonymous):

sure looks good to me!

OpenStudy (anonymous):

that y'=(-5-6x-y)/x

OpenStudy (anonymous):

but how do i finish it?

OpenStudy (amistre64):

by inserting whatever point they give you to fill in x and y

OpenStudy (anonymous):

you have a curve and have decided that y is a (implicit) function of x. since you did not solve for y in terms of x your answer must have a y in it. you are done

OpenStudy (anonymous):

haha there is no point i wish there was

OpenStudy (amistre64):

then y' = what it equals and you are done with it

OpenStudy (anonymous):

o ok..so i had the answer right then

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

ok thanks